An aluminum ball X and an iron ball Y of the same volume are thrown horizontally with the same velocity from the top of a building. Neglecting air resistance, X reaches the ground (a) before Y and at the same distance from the building (b) at the same time as Y and at a nearer distance to the building (c) at the same time as Y and at the same distance from the building (d) at the same time as Y and further from the building (e) after Y and at the same distance from the building
1
Expert's answer
2020-11-16T07:54:05-0500

Explanations & Calculations


  • Since the density of Iron is greater than that of Aluminum, mass of the ball Y is greater than X's.

"\qquad\qquad\small m=V\times \rho"

  • When they are thrown, they move in projectiles under the gravitational acceleration which is constant throughout the flight of the 2 balls.
  • If you closely inspect the 4 equations of motion (these are applicable only for a constant acceleration), "\qquad\qquad
\begin{aligned}
\small v&=\small u+at\\
\small v^2&=\small u^2+2as\\
\small s&=\small ut+\frac{1}{2}at^2\\
\small s&=\small \frac{(v+u)t}{2}
\end{aligned}" you will see that there is no any influence of the mass of an object over the nature of the flight. Only what happens is "\small a \to g".



  • Therefore, when the influence of air is neglected (in reality we experience the air resistance) both the balls experience the same flight giving (C) as the answer.



  • If the height of the building is h & it takes time t for the ball to reach the ground & apply "\small s=ut+\frac{1}{2}at^2" downward. Then, "\small h=0t_1+\frac{1}{2}gt_1^2 \to t_1=\sqrt{\frac{2h}{g}}"
  • Therefore, "\small \bold{t_X=t_Y}"


  • If a ball travels some s1 distance horizontally during that time, apply "\small s= ut+\frac{1}{2}at^2" for the horizontal motion. Then, "\small s_1 =u_0t_1+\frac{1}{2}\times 0\times t_1\to s_1=u_0t_1 \to s_1=u_0\sqrt{\frac{2h}{g}}"
  • Therefore, "\small \bold{s_X=s_Y}"